# Offline Model-Based Optimization: Comprehensive Review

### v1

@@ 6 Discussion and Future Direction
Robust and Realistic Benchmarking
Current benchmarks in offline MBO face two major challenges. First, some benchmarks—such as TFB8 and TFB10 (Barrera et al., 2016a)—offer overly constrained search spaces where even simple gradient ascent methods can achieve impressive results, making it difficult to distinguish the performance of more sophisticated algorithms. Second, benchmarks like superconductor (Hamidieh, 2018) often rely on learned oracles for evaluation, which can be vulnerable to manipulation and may not accurately reflect true performance. Moving forward, it is essential to develop benchmarks that not only present more challenging search spaces but also incorporate rigorous, reliable evaluation protocols that are resistant to exploitation.
@@
In offline optimization, capturing the significant epistemic uncertainty in surrogate models is paramount, especially because identifying which inputs drive this uncertainty remains a core challenge. While uncertainty estimation is similarly important in online optimization, in offline scenarios it is even more critical for preventing reward hacking, as one cannot correct for flawed model predictions through active data collection. Various techniques—such as adversarial regularization (Trabucco et al., 2021), smoothness priors (Yu et al., 2021), and kernel-based methods (Chen et al., 2022b)—have been proposed to mitigate uncertainty and ensure safer optimization; however, there has been comparatively little focus on leveraging Bayesian methods for directly quantifying epistemic uncertainty. Moreover, existing benchmarks often emphasize overall optimization performance without clarifying whether observed gains stem from superior surrogate modeling, improved optimization strategies, or mere chance. This lack of distinction underscores the need for independent and rigorous evaluations of the uncertainty estimation capabilities of the surrogate model in newly developed algorithms. Although recent efforts—such as Jain et al. (2022) using Monte Carlo (MC) dropout (Gal & Ghahramani, 2016) and deep ensembles (Lakshminarayanan et al., 2017), and Kim et al. (2024a) incorporating uncertainty measures for conservative search—have made inroads into this area, they still fall short of providing a robust solution. Consequently, advancing offline MBO demands a fundamentally stronger approach to uncertainty quantification that transcends basic MC dropout or ensembling techniques.
@@
  • DEUP: Direct Epistemic Uncertainty Prediction (Lahlou et al., 2023): This approach directly estimates the excess risk associated with model misspecification by learning a secondary predictor for generalization error, offering a more principled uncertainty measure.
@@
  • Deep Kernels: Deep kernels, which leverage deep architectures to learn expressive kernel functions, offer a scalable alternative to Gaussian processes (Wilson et al., 2016; Liu et al., 2020). They have been applied to biological sequence design scenarios by employing a denoising autoencoder to learn a discriminative deep kernel GP for Bayesian optimization of biological sequences (Stanton et al., 2022). This approach was further refined by Gruver et al. (2023), who used a discrete diffusion model and simple last-layer ensemble techniques to evaluate uncertainty. These successes in online optimization could be directly transferred to offline MBO, where uncertainty estimation helps prevent reward hacking. Such a strategy could effectively combine existing surrogate modeling and generative modeling methods in offline MBO fields.
@@
Offline surrogate models often suffer from limited data coverage, making them vulnerable to overestimation and reward hacking in regions outside the training distribution. A promising mitigation strategy is to leverage factorized graphical models—such as Functional Graphical Models (FGMs) (Kuba et al., 2024b)—to decompose a high-dimensional function into a sum of local subfunctions defined over smaller cliques. This structured factorization enables each subcomponent to be learned from a subspace with denser data coverage, thereby localizing distribution shifts and enhancing both robustness and OOD generalization (Kuba et al., 2024a). In addition to FGMs, candidate methods for graph discovery – such as those based on GFlowNets –offer alternative avenues to uncover latent causal structures (Deleu et al., 2022). Employing such graphical discovery methods will provide a principled means to address distributional shifts in offline MBO, ultimately leading to more reliable optimization outcomes under limited data coverage.
@@
While offline MBO has traditionally been applied to design tasks in fields such as biology or chemistry, its methodologies can also be used for post-training large language models (LLMs). Approaches like supervised fine-tuning (SFT), RLHF (Ouyang et al., 2022), or RL-based reasoning (Guo et al., 2025), aim to enhance text generation by leveraging reward signals derived from human preference models or other evaluative metrics. Because these reward models are inherently uncertain and prone to reward hacking, adopting the conservative and safe optimization techniques developed in offline MBO can help mitigate such risks. Moreover, since reward hacking can lead to catastrophic misalignment of superhuman-level intelligence toward harmful ends (Bengio et al., 2025), integrating robust optimization algorithms from the offline MBO community can further contribute to LLM safety.
@@ Acknowledgement
This research is partially supported by the Fonds de recherche du Québec – Nature et technologies. This research also is partially supported by funding from Samsung, CIFAR and the CIFAR AI Chair program. We sincerely thank Moksh Jain from Mila for the insightful discussions.

### v2

@@ 6 Discussion and Future Direction
Robust and Realistic Benchmarking on Scientific Design
Current scientific design benchmarks in offline MBO face two major challenges. First, due to the difficulty of collecting labeled data, some benchmarks—such as TFB8 and TFB10 (Barrera et al., 2016)—offer overly constrained search spaces where even simple gradient ascent methods can achieve impressive results, making it difficult to distinguish the performance of more sophisticated algorithms. Second, benchmarks like superconductor (Hamidieh, 2018) often rely on learned oracles for evaluation, which can be vulnerable to manipulation and may not accurately reflect true performance. Moving forward, it is essential to develop benchmarks that not only present more challenging search spaces but also incorporate reliable evaluation protocols that are resistant to exploitation.
@@
In offline black-box optimization, capturing the significant epistemic uncertainty in surrogate models is paramount, especially because identifying which inputs drive this uncertainty remains a core challenge. While uncertainty estimation is similarly important in online black-box optimization, in offline scenarios it is even more critical for preventing reward hacking, as one cannot correct for flawed model predictions through active data collection. Various techniques—such as adversarial regularization (Trabucco et al., 2021), smoothness priors (Yu et al., 2021a), and kernel-based methods (Chen et al., 2022b)—have been proposed to mitigate uncertainty and ensure safer optimization; however, there has been comparatively little focus on leveraging Bayesian methods for directly quantifying epistemic uncertainty. Moreover, existing benchmarks often emphasize overall optimization performance without clarifying whether observed gains stem from superior surrogate modeling, improved optimization strategies, or mere chance. This lack of distinction underscores the need for independent and rigorous evaluations of the uncertainty estimation capabilities of the surrogate model in newly developed algorithms. Although recent efforts—such as Jain et al. (2022) using Monte Carlo (MC) dropout (Gal & Ghahramani, 2016) and deep ensembles (Lakshminarayanan et al., 2017), and Kim et al. (2024a) incorporating uncertainty measures for conservative search—have made inroads into this area, they still fall short of providing a robust solution. Consequently, advancing offline MBO demands a fundamentally stronger approach to uncertainty quantification that transcends basic MC dropout or ensembling techniques.
@@
  • DEUP: Direct Epistemic Uncertainty Prediction (Lahlou et al., 2023b): This approach directly estimates the excess risk associated with model misspecification by learning a secondary predictor for generalization error, offering a more principled uncertainty measure.
@@
  • Deep Kernels: Deep kernels, which leverage deep architectures to learn expressive kernel functions, offer a scalable alternative to Gaussian processes (Wilson et al., 2016; Liu et al., 2020). They have been applied to biological sequence design scenarios by employing a denoising autoencoder to learn a discriminative deep kernel GP for Bayesian optimization of biological sequences (Stanton et al., 2022). This approach was further refined by Gruver et al. (2023), who used a discrete diffusion model and simple last-layer ensemble techniques to evaluate uncertainty. These successes in online black-box optimization could be directly transferred to offline MBO, where uncertainty estimation helps prevent reward hacking. Such a strategy could effectively combine existing surrogate modeling and generative modeling methods in offline MBO fields.
@@
Offline surrogate models often suffer from limited data coverage, making them vulnerable to overestimation and reward hacking in regions outside the offline distribution. A promising mitigation strategy is to leverage factorized graphical models—such as Functional Graphical Models (FGMs) (Kuba et al., 2024b)—to decompose a high-dimensional function into a sum of local subfunctions defined over smaller cliques. This structured factorization enables each subcomponent to be learned from a subspace with denser data coverage, thereby localizing distribution shifts and enhancing both robustness and OOD generalization (Kuba et al., 2024a). In addition to FGMs, candidate methods for graph discovery – such as those based on GFlowNets –offer alternative avenues to uncover latent causal structures (Deleu et al., 2022). Employing such graphical discovery methods will provide a principled means to address distributional shifts in offline MBO, ultimately leading to more reliable optimization outcomes under limited data coverage.
@@
While offline MBO has traditionally been applied to design tasks in fields such as biology or chemistry, its methodologies can also potentially inform post-training strategies for large language models (LLMs). Approaches like supervised fine-tuning (SFT), RLHF (Ouyang et al., 2022), or RL-based reasoning (Guo et al., 2025), aim to enhance text generation by leveraging reward signals derived from human preference models or other evaluative metrics. Because these reward models are inherently uncertain and prone to reward hacking, adopting the conservative and safe optimization techniques developed in offline MBO may help mitigate such risks. Looking forward, we speculate that insights from offline MBO’s robust optimization principles may also inform future directions in ensuring the safety and reliability of increasingly capable LLMs (Bengio et al., 2025).
@@ Acknowledgement
This research is partially supported by the Fonds de recherche du Québec – Nature et technologies, as well as funding from Samsung, CIFAR, the CIFAR AI Chair program, and the Canadian AI Safety Institute Research Program at CIFAR through a Catalyst award. Minsu Kim acknowledges funding from the KAIST Jang Young Sil Fellow Program. We sincerely thank Moksh Jain from Mila and Brandon Trabucco from CMU for insightful discussions.